OpenStudy (anonymous):

The number of industrial accidents this month fell to 963 accidents from 1239 accidents last month. Find the percent of decrease. ( round to the nearest tenth as needed)

OpenStudy (anonymous):

i did the same formula that @agent0smith told me but i got it wrong

OpenStudy (anonymous):

so it would be % of change = 1239-963/1239 x 100%

OpenStudy (agent0smith):

Hopefully you used parentheses...

OpenStudy (anonymous):

i did !

OpenStudy (agent0smith):

You may have to put the percentage in as negative. It should be (new-old)/old... you have (old-new)/old

OpenStudy (anonymous):

what do you mean?

OpenStudy (calculusfunctions):

% change = [(new quantity - old quantity)/(old quantity)] X 100 %

OpenStudy (calculusfunctions):

If this change is negative, then there is a decrease.

OpenStudy (calculusxy):

Lets use a ratio first: Percent increase/100=Amount of increase/Original Quantity

OpenStudy (anonymous):

so can you help me plug in the number and show me the steps

OpenStudy (calculusxy):

Sure

OpenStudy (calculusxy):

Okay so first we need to solve for the percent decrease. To do that, we have to substitute for the numbers in the right side of the equal sign.

OpenStudy (agent0smith):

Use the formula i and calculus functions gave % change = [(new quantity - old quantity)/(old quantity)] X 100 % new quantity = 963 old quantity = 1239

OpenStudy (calculusxy):

Sorry it wasn't "Percent Increase" but "Percent Decrease"

OpenStudy (anonymous):

okay its fine

OpenStudy (calculusxy):

To compute the Amount of Decrease, we have to subtract 1239-963=276. To find out the original amount, we know that the amount that we are changing from is 1239 and moving onto 963. Since the new quantity is 963 and the old quantity is 1239, thats what you will get off with. This will make the right side of the equation as 276/1239. % of Increase/100=276/1239

OpenStudy (agent0smith):

@cmenendez55 the only mistake you made was the numbers the wrong way around: (1239-963)/1239 x 100% should be (963-1239)/1239 x 100%

OpenStudy (anonymous):

0.0496 but rounded is 0.050 right?

OpenStudy (calculusxy):